I have been in geometry for three days and my teacher has given us one example and i just don't know how to even start to do them. we don't even have a text book. if you could help me i would really appreciate it. the problem is Give a line segment AD with points B and C between A and D. If AB=CD, then AC=BD.
Answer by josmiceli(19441)

The line segment is
a-b-c-d
You are given that ab = cd
from looking at the line I see that
(1) ac = ab + bc
I can also see that
(2) bd = bc + cd
---------------------
Now you are told to show that
ac = bd
---------------------
From (1),
ab = ac - bc
From (2)
cd = bd - bc
--------------------
If ab = cd as given,
ac - bc = bd - bc
Add bc to both sides and you get
ac = bd